# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

### v2

@@ Abstract
Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. This paper introduces an integrated framework that leverages the capabilities of the Transformer model-based neural networks and deep reinforcement learning (DRL) algorithms to optimize system maintenance actions. Our approach employs the Transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the remaining useful life (RUL) of an equipment. Additionally, the DRL component of our framework provides cost-effective and timely maintenance recommendations. We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions, compared to the other prevalent machine learning-based methods. Our proposed approach provides an innovative data-driven framework for industry machine systems, accurately forecasting equipment lifespans and optimizing maintenance schedules, thereby reducing downtime and cutting costs.
@@ I Introduction
The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies [1]. With these advancements, however, comes the increased requests for reliable, efficient, and cost-effective maintenance systems. Prescriptive maintenance has thus emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [2]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [3], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning. However, current predictive maintenance systems largely rely on heuristic models for action planning or conventional machine learning algorithms for sensor data processing. While effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues like vanishing or exploding gradients, limiting their ability to accurately predict Remaining Useful Life (RUL).
@@
Apart from that, there is a trend in recent works to leverage Reinforcement Learning (RL) to make plans in maintenance tasks [4]. RL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks. As a result, the RL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs. However, these methods consistently encounter the cold start problem [5], coupled with low learning efficiency. In some special cases such as aeroengine, these methods also suffer from a lack of reliability and explainability. This often leads to their decisions being perceived as untrustworthy by users.
@@
Unlike existing solutions that address either predictive or prescriptive maintenance, we propose a framework named TranDRL, which integrates both aspects seamlessly. TranDRL leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations. By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges by Transformer neural networks and Reinforcement Learning with Human Feedback (RLHF) algorithms. Transformers excel at modeling long-term dependencies in sequential data by the attention mechanism and are highly adaptable, making them ideal for handling complex industrial sensor data. By leveraging the expertise and insights of human operators or maintenance professionals, the algorithm can efficiently guide its actions, accelerating the convergence of the DRL algorithm towards optimal maintenance strategies. This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and ultimately leading to more efficient and cost-effective industrial maintenance practices.
Contributions. The contributions of this paper can be summarized as follows:
@@
  • To improve the accuracy of the RUL prediction module, we adopt Transformer to address the difficulty in reflecting long-term dependencies of sequential sensor data in industrial scenarios.
@@
  • We propose a DRL-based recommendation system for determining maintenance actions, reducing both labor involvement and operational costs. Furthermore, we utilize RLHF to confirm the correct maintenance actions and improve learning efficiency.
  • A Transformer-DRL framework is proposed to optimize the prescriptive maintenance process, which can be extended to multiple manufacturing scenarios as well.
@@
  • Extensive simulation experiment results, conducted on different datasets, demonstrate that our method outperforms baseline methods in terms of both predicting RUL and optimizing maintenance actions.
@@
Organizations. The rest of the paper is organized as follows. Section II reviews related work in the field of prescriptive maintenance utilizing machine learning algorithms. Section III formulate the problem we are trying to solve. Then, section IV describes our proposed framework for Transformer-driven prediction maintenance in detail. Next, section V presents a case study on our proposed framework using a real-world dataset. Finally, section VI concludes the paper and outlines future directions for research. Notations used across the paper is provided in Table I.
TABLE I: Notations.
Symbol Description
$\mathbf{X}$
Time series of sensor readings and operational parameters.
$R(t)$
Remaining Useful Life at time $t$: $R(t)=T-t$.
$T$
Total expected lifespan of the component.
$t$
Current operational time of the machine.
$M$
Model estimating $R(t)$ from $\mathbf{X}$.
$s$
State encapsulating health and RUL metrics.
$\mathcal{A}$
Set of maintenance actions.
$a$
Action chosen to maximize future reward.
$a^{*}$
Optimal action maximizing cumulative benefit.
$Q(s,a)$
Expected reward of action $a$ in state $s$.
$\epsilon$
Exploration rate in DRL.
$\alpha$
Learning rate in DRL.
$\gamma$
Discount factor for future rewards.
$\bm{y}$
Predicted RUL output by Transformer model.
$\bm{s}$
Observation state in DRL from RUL predictions.
$d_{\text{model}}$
Input/output dimension in Transformer.
$d_{k},d_{v}$
Dimensions of keys/queries and values in attention.
$\mathbf{Q},\mathbf{K},\mathbf{V}$
Query, key, and value matrices in attention.
$FFN(\cdot)$
Feed-forward network in Transformer.
$\bm{\Lambda}$
Attention weight matrix.
$\alpha_{ij}$
Attention weight from $j$-th input to $i$-th output.
$\bm{h}$
Hidden states in encoder.
O
Output states from decoder.
$\sigma(\cdot)$
Activation function in prediction module.
$K$
Number of machines in federated learning setup.
$f_{\text{ensemble}}$
Function for RUL prediction in multi-machine case.
$\mathbf{W},b$
Weight matrix and bias in prediction module.
PCA
Used for selecting important parts of RUL.
@@
II Related Work
[…]
Corrective maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected. Preventive maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs. These two types of maintenance methods are traditionally the most commonly adopted in the industry. However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance. Predictive maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as condition-based maintenance. Prescriptive maintenance is generally considered an extension of predictive maintenance, optimizing maintenance strategies based on fault diagnosis results. It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [7]. Compared to traditional maintenance methods, prescriptive maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
II-A Deep Learning-Based Remaining Useful Life Prediction
In production maintenance tasks, accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [8, 9]. Notably, they also entail labor-intensive manual feature engineering.
However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, boosted by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [10, 11]. Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language processing tasks [12], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [13] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
II-B Deep Reinforcement Learning enabled Predictive Maintenance
DRL’s concept into predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 4]. While these applications exhibit DRL’s capability in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, focusing on actionable preventative measures. Gordon et al. [15] explore into employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
In existing maintenance studies, we have explored a lot at preventive and predictive methods, as mentioned in [16] and [17]. But now, there is growing interest in prescriptive maintenance, which goes beyond just predicting problems to suggesting the best actions to take. Even though the work in [16] and [17] has started to light the way, most efforts are either focused on predicting what will happen or on prescribing what to do. Even with improvements in using Transformer technology for predictions and DRL for making recommendations, there is no plan to integrate these two ideas together. Our study aims to fill this gap by introducing a way to use both the forecasting power of Transformers and the decision-making abilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
II-C Deep Reinforcement Learning with Human Feedback
Inspired by preference theory [18], Reinforcement Learning from Human Feedback (RLHF) improves the performance of DRL by leveraging feedbacks from human, and releases the difficulty of designing reward function by incorporating human guidance and expertise [19]. To tackle the problem of informative limitation and expensive time consumption, Ibarz et al.  [20] combine expert demonstrations and trajectory preferences together and attempt to let human interact with agent directly. To enhance sample efficiency, other follow-up works leverage semi-supervised learning and data augmentation techniques into RLHF [21, 22]. Most previous works are under the non-Markov assumption which poses a threat to modeling reward function, Early et al.  [23] release this assumption and formulate the scenario into a multiple instance learning process. With the prosperity of large language model recently, RLHF also shows its power in fine-tuning step before the model development in this specific domain [24, 25]. Furthermore, in safety-critical applications like aeroengine maintenance, deviating from standard procedures can have grave consequences. Such non-standard maintenance practices might result in serious safety incidents. Combining human feedback can also help in setting safe boundaries and ensuring that the agent’s behavior remains within acceptable limits. To make the policy more reliable, the incorporation of human feedback can make the agent’s behavior more predictable and better aligned with human expectations, enhancing the explainability of its actions and building trust among users.
III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. However, with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it also affects aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [26]. Therefore, the primary challenge is to recommend proper maintenance actions effectively based on accurate prediction of failures of the machine.
III-A Predicting Remaining Useful Life (RUL)
Predicting RUL is essential for predictive maintenance, which provides an estimation of the duration that a machine or its components will continue to function before encountering a failure. An accurate RUL prediction enable the industries to strategically schedule their maintenance activities. The RUL offers a dynamic and predictive window into the operational health of a machine. The operational data of a machinery component is typically represented as a time series $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $t$ represents the current age or operational time of the machine.
The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors, for example, machine conditions. The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $\mathbf{X}$ to provide an accurate approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
III-B Optimal Maintenance Action Recommendations with DRL
While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions. Upon gaining prediction results of potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
The essence of this problem is rooted in the principle of maximizing long-term rewards. We denote a value function $Q(s,a)$, which characterizes the expected cumulative benefit or reward of adopting action $a$ in state $s$. The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
  • $Q(s,a)$ encompasses immediate benefits and projected future rewards, taking into consideration the dynamic nature of machinery health and operation.
  • The cumulative reward should integrate both current and weighted future’s benefits strategically.
@@ IV Proposed Framework
To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions as illustrated in Figure 1 and Algorithm 1. Specifically, we consider two cases for the RUL prediction, including a single machine case and a multi-machine case.

### v3

@@ Abstract
Industrial systems require reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. Existing studies rely on heuristic models which may struggle to capture complex temporal dependencies. This paper introduces an integrated framework that leverages the capabilities of the Transformer and Deep Reinforcement Learning (DRL) algorithms to optimize system maintenance actions. Our approach employs the Transformer model to effectively capture complex temporal patterns in IoT sensor data, thus accurately predicting the Remaining Useful Life (RUL) of equipment. Additionally, the DRL component of our framework provides cost-effective and timely maintenance recommendations. Numerous experiments conducted on the NASA C-MPASS dataset demonstrate that our approach has a performance similar to the ground-truth results and could be obviously better than the baseline methods in terms of RUL prediction accuracy as the time cycle increases. Additionally, experimental results demonstrate the effectiveness of optimizing maintenance actions.
@@ I Introduction
Industry 4.0 has introduced unprecedented levels of automation and data exchange in manufacturing technologies [1]. However, with these advancements comes increased requirement for reliable, efficient, and cost-effective maintenance systems. Prescriptive maintenance has thus emerged as a crucial topic in both research and industrial sectors, offering the promise of proactive fault detection and optimized operation strategy planning [2]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [3], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning. However, current predictive maintenance systems largely rely on heuristic models for action planning or conventional machine learning algorithms for sensor data processing. Although effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues such as vanishing or exploding gradients, limiting their ability to accurately predict RUL. To address these issues, a number of new frameworks (e.g., attention in recurrent networks [4]) are developed to capture the correlation in temporal sequence data.
@@
Apart from that, there is a trend in recent studies to leverage DRL to make plans in maintenance tasks [5]. DRL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks. As a result, the DRL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs. However, these methods consistently encounter the cold start problem [6], coupled with low learning efficiency. In some special cases, such as aero-engines, these methods also suffer from a lack of reliability and explainability. This often leads to their decisions being perceived as untrustworthy by users.
@@
Unlike existing solutions, we propose a framework named TranDRL, which integrates both aspects seamlessly. TranDRL leverages both the Transformer framework and DRL to not only accurately predict the RUL but also convert these predictions into maintenance action recommendations. Additionally, we adopt the RL with Human Feedback (RLHF) algorithm, which incorporates the human expert’s opinion on maintenance actions into the DRL algorithm to facilitate the algorithm in making decisions. By integrating these two critical neural networks and human expert’s experience, our framework provides a comprehensive solution for industrial maintenance challenges. Transformers excel at modeling long-term dependencies in sequential data through their attention mechanism and exhibit a high degree of adaptability, making them ideal for handling complex industrial sensor data. Utilizing the expertise and insights of human maintenance professionals, the algorithm can efficiently guide the production operational actions, expediting the convergence of the DRL algorithm toward optimal maintenance strategies. This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and leading to more efficient and cost-effective industrial maintenance practices.
Contributions. The main contributions of this paper can be summarized as follows:
@@
  • To improve the accuracy of the RUL prediction module, we adopt a Transformer-based neural network to address the difficulty in reflecting long-term dependencies of sequential sensor data.
@@
  • We propose a DRL-based recommendation approach for determining maintenance actions, reducing both labor involvement and operational costs. Furthermore, we utilize the RLHF algorithm to confirm the correct maintenance actions and improve the learning efficiency of the DRL framework.
  • The transformer-enabled DRL algorithm is tailored into a federated learning framework, which can optimize the prescriptive maintenance process in multi-machine scenarios.
@@
  • Extensive simulation experiment results demonstrate that our method outperforms a series of baseline methods.
@@
Organizations. The rest of the paper is organized as follows. Section II reviews related work in the field of prescriptive maintenance utilizing machine learning algorithms. Section III formulates the problem we are trying to solve. Then, Section IV describes our proposed framework for Transformer-driven prediction maintenance in detail. Next, Section V presents a case study on our proposed framework using a real-world dataset. Finally, Section VI concludes the paper and outlines future directions for research. Notations used across the paper are provided in Table I.
TABLE I: Notations and abbreviations.
Symbol Descriptions
$\mathbf{X}$ Time series of sensor readings and operational parameters.
$R(t)$ Remaining Useful Life at time $t$: $R(t)=T-t$.
$T$ Total expected lifespan of the component/machine.
$t$ Current operational time of the machine in concern.
$M$ Model estimating $R(t)$ from $\mathbf{X}$.
$s$ State encapsulating the health and RUL metrics.
$\mathcal{A}$ Set of maintenance actions.
$a$ Action chosen to maximize the future reward.
$a^{*}$ Optimal action maximizing the cumulative benefit.
$Q(s,a)$ Expected reward of action $a$ in state $s$.
$\epsilon$ Exploration rate in DRL.
$\alpha$ Learning rate in DRL.
$\gamma$ Discount factor for future rewards.
$\bm{y}$ Predicted RUL output by the neural network model.
$\bm{s}$ Observation state in DRL from RUL predictions.
$d_{\text{model}}$ Input/output dimension of the Transformer.
$d_{k},d_{v}$ Dimensions of keys/queries and values in attention computation.
$\mathbf{Q},\mathbf{K},\mathbf{V}$ Query, key, and value matrices in attention computation.
$FFN(\cdot)$ Feed-forward network in the Transformer module.
$\bm{\Lambda}$ Attention weight matrix.
$\alpha_{ij}$ Attention weight from the $j$-th input to the $i$-th output.
$\bm{h}$ Hidden states in the encoder.
O Output states from the decoder.
$\sigma(\cdot)$ Activation function in the prediction module.
$K$ Number of machines in federated learning setups.
$f_{\text{ensemble}}$ Function for RUL prediction in the multi-machine case.
$\mathbf{W},b$ Weight matrix and bias in the prediction module.
PCA Principal component analysis.
DQN Deep Q-network.
PPO Proximal Policy Optimization.
SAC Soft Actor-Critic.
II Related Work
Prescriptive maintenance uses historical operational data and (possibly implicit) system models to optimize maintenance timing and strategies. It extends the prediction maintenance approaches by optimizing strategies based on fault diagnosis, focusing on identifying root causes and selecting proper maintenance actions [7]. Compared to traditional methods, it aims to prevent waste and production losses by leveraging historical system status to forecast future machine status and optimize maintenance resources.
II-A Deep Learning-based Remaining Useful Life Prediction
In production maintenance tasks, accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often require deep domain-specific knowledge, especially for tasks such as signal denoising and feature extraction [8, 9]. Notably, they also entail labor-intensive manual feature engineering. However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, boosted by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained popularity owning to their proficiency in fault diagnosis and predictive maintenance [10, 11]. Their intrinsic capability to distill hierarchical features from raw data eliminates the need of manual feature engineering, thus fostering more accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language processing tasks [12], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [13] integrated a time-embedding layer, augmenting the Transformer’s time series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
II-B Deep Reinforcement Learning-enabled Predictive Maintenance
DRL’s concept of predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 5]. While these applications exhibit DRL’s capability in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, focusing on actionable preventative measures. Recent work such as [15] explored the possibility of employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
In existing maintenance studies, researchers have explored a number of preventive and predictive methods, as mentioned in [16] and [17]. However, there is a growing interest in prescriptive maintenance, which extends beyond merely predicting problems, instead offering suggestions for the best actions to take. Studies such as [16] and [17] have started to light the way, with most efforts focused on either predicting what will happen or prescribing what to do. However, even with improvements in using the Transformer technology for predictions and DRL for making recommendations, there is yet no plan to integrate these two ideas. Our study aims to fill this gap by introducing a way to utilize both the forecasting power of Transformers and the decision-making capabilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
Furthermore, inspired by preference theory [18], RLHF improves DRL performance by leveraging human feedback, reducing the difficulty in designing reward functions [19]. To address limitations of information and expensive data collection, expert demonstration and trajectory preferences are combined in [20], enabling direct human-agent interaction. Subsequent studies enhanced sample efficiency through semi-supervised learning and data augmentation [21, 22]. In [23] the non-Markov assumption is relaxed, formulating the scenario as a multiple-instance learning process. In safety-critical applications such as aero-engine maintenance, human feedback can set safe boundaries, therfore ensuring that the behavior of an agent remains within acceptable limits, making it more predictable and aligned with human expectations, as well as enhancing explainability and trust.
III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. However, with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it also affects aspects such as increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [24]. Therefore, the primary challenge is to recommend maintenance actions based on accurate machine failure prediction effectively.
III-A Remaining Useful Life Prediction
Predicting RUL is essential for predictive maintenance, which provides an estimation of the duration that a machine or its components will continue to function before encountering a failure. An accurate RUL prediction enables industries to strategically schedule their maintenance activities. The RUL offers a dynamic and predictive window into the operational health of a machine. The operational data of a machinery component is typically represented as a time series $\mathbf{X}=[\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}}]$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions at a sampling instance $i$. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $t$ represents the machine’s current age or operational time.
The inherent complexity is that $T$ is not a constant but a variable that might change based on multiple factors such as machine conditions. The primary challenge is to develop a predictive model $M$ that can use the time-series data $\mathbf{X}$ to provide an approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
III-B Optimal Maintenance Action Recommendation
While predicting the RUL can help industries predict potential downtimes, it is the actionable insights derived from these predictions that empower industries to make informed decisions. Upon gaining prediction results of potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to extend its working time and ensure optimal performance.
Therefore, the challenge of this problem is to maximize delayed reinforcements. We both present and denote a value function $Q(s,a)$ that expresses the sum of the utility or profit from choosing $a$ in $s$. The objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
  • $Q(s,a)$ reflects immediate benefits and projected future rewards, considering the dynamics of machinery health and operation.
  • The cumulative reward should integrate the current and the weighted future benefits.
@@ IV Proposed Framework
To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions, as illustrated in Algorithm 1 and Figure 1. Specifically, the sensor data is collected and utilized to predict the RUL of machines, after which DRL algorithms are employed to determine the optimal action for repairing machines.